### v1

# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

Abstract
Extending the context window in large language models (LLMs) is essential for applications involving long-form content generation. However, the linear increase in key-value (KV) cache memory requirements and the quadratic complexity of self-attention with respect to sequence length present significant challenges during fine-tuning and inference. Existing methods suffer from performance degradation when extending to longer contexts. In this work, we introduce a novel context extension method that optimizes both fine-tuning and inference efficiency. Our method exploits a key observation: in the frequency domain, the energy distribution of the KV cache is primarily concentrated in low-frequency components. By filtering out the high-frequency components, the KV cache can be effectively compressed with minimal information loss. Building on this insight, we propose an efficient compression technique, FreqKV, that iteratively compresses the increasing KV cache to a fixed size in the frequency domain, applicable to both fine-tuning and inference. FreqKV introduces no additional parameters or architectural modifications. With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window efficiently. Experiments on various long context language modeling and understanding tasks demonstrate the efficiency and efficacy of the proposed method.
1 Introduction
Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process. However, it is inevitable for LLMs to process sequences that exceed the preset context size. LLMs struggle to maintain their performance when generalized to longer contexts. Additionally, the computation cost of the self-attention mechanism [23] grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference. They evict [26, 17] or merge [29, 24] KV states of less important tokens following certain rules. They use attention scores to measure the importance and approximate the original full attention. However, while these methods provide an approximation of full computation on existing tokens through different strategies, they can not fully prevent performance degradation when decoding future tokens.
Recent studies propose to fine-tune LLMs to longer contexts to extend the context window. LongLoRA [7] trains LLMs using shifted sparse attention. Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence. Concurrently, LoCoCo [5] and Activation Beacon [28] introduce additional modules to compress KV states. They incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) [27]. Moreover, Fourier Transformer [13] discards the high-frequency parts of the contexts and downsample the hidden states in the encoders. Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs. We transform key states and value states in LLaMA-2-7b [22] from the time domain to the frequency domain for power spectrum analysis. As shown in Figure 1, the energy distribution increasingly concentrates on low-frequency components as the computation process progresses along the decoder layers. It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
In this paper, we introduce FreqKV, an efficient context extension method that iteratively compresses key-value states in the frequency domain. Compression is triggered only when the KV cache reaches the predefined context window size. During each compression step, the low-frequency components of the KV states are preserved at a specified retaining ratio. Subsequent tokens are appended to the compressed cache until it is filled again. This ensures that the cached KV states that each query token can attend to is limited below the context window size. To reduce memory and computational costs, the compressed cache will be further compressed together with the incoming tokens. This iterative compression mechanism leads to an increased compression level of the earlier contexts as the sequence length grows. Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts. FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling. Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long-context understanding.
2 Related Work
KV Compression for LLMs.
To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model. One common approach is selective token eviction [26, 17, 6], where less significant tokens are discarded. Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens. To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts [29, 24, 25]. However, these inference methods often sacrifice performance for efficiency.
Context Extension for LLMs.
Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs. LongLoRA [7] employs shifted sparse attention during the parameter-efficient fine-tuning. However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training. Other techniques, such as LoCoCo [5], integrate convolutional operations into LLMs for compressing long contexts. They fine-tune the compression modules together with LLMs. Landmark attention [19] uses landmark tokens to retrieve previous input blocks. Similarly, Activation Beacon [28] introduces a special token to represent the previous context for compression. However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models. In contrast, our proposed method achieves context extension without introducing any additional parameters.
Learning in the Frequency Domain.
Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs [11]. It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels [27]. These works have inspired efforts to process natural language. FNet [16] enhances the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens. Additionally, Fourier Transformer [13] eliminates redundancies in the context through frequency domain processing within encoder architectures.
However, because of the auto-regressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs. To the best of our knowledge, FreqKV is the first work that explores compressing key-value states in the frequency domain for decoder-only LLMs.
3 Preliminaries
3.1 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain. Several variants of the DCT exist, with DCT-II being the most common. For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
$y_{t}=\alpha_{t}\sum_{n=0}^{N-1}x_{n}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right],$ $\alpha_{t}=\begin{cases}\sqrt{\frac{1}{N}}&\text{ if }t=0,\\ \sqrt{\frac{2}{N}}&\text{ otherwise }\end{cases}$ (1)
where $t=0,1,\cdots,N-1$. $\alpha_{t}$ is the normalization factor.
The time-domain signal ${\bm{X}}_{0:N-1}$ can be recovered by applying the inverse DCT (IDCT) on the frequency components ${\bm{Y}}_{0:N-1}$:
$x_{n}=\sum_{t=0}^{N-1}\alpha_{t}\cdot y_{t}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right].$ (2)
The frequency components are expressed as a combination of the original signals. The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$. The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components. The components of higher energy in the frequency domain indicate they are more informative [13].
3.2 Self-Attention
For the incoming token $x_{N}$, the prefilled $N$ tokens $X_{0:N-1}$ are utilized as the cache during decoding. Denote the hidden states of the $N+1$ tokens input to a specific layer of LLMs as ${\bm{H}}_{0:N}=[\displaystyle{\bm{h}}_{0},\cdots,\displaystyle{\bm{h}}_{N}]$. The query, key and value states of ${\bm{x}}_{N}$ are computed as follows::
${\bm{q}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{Q},\quad\displaystyle{\bm{k}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{v}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{V},$ (3)
where ${\bm{W}}^{Q},\displaystyle{\bm{W}}^{K},\displaystyle{\bm{W}}^{V}$ are the projection matrices for the query, key and value states, respectively. For simplicity, indices corresponding to layers and heads have been omitted.
The cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ are:
${\bm{K}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{V}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{V}.$ (4)
When calculating attention scores, the incoming token $x_{N}$ attends to all cached KV states as well as to itself:
$\mathcal{\displaystyle{\bm{A}}}(N)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{N}[\displaystyle{\bm{K}}_{0:N-1}\oplus\displaystyle{\bm{k}}_{N}]^{T}}{\sqrt{d}}\right)\cdot[\displaystyle{\bm{V}}_{0:N-1}\oplus\displaystyle{\bm{v}}_{N}],$ (5)
where $d$ is the hidden dimension. $\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$
4 FreqKV
(a) The average power spectrums of key states.
(b) The average power spectrums of value states.
Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b. 1000 documents are sampled from CNN/Daily Mail [14]. We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
4.1 Energy Concentration in the Frequency Domain
We transform key states and value states from the time domain, which is the sequence dimension, to the frequency domain. The average power spectrums in different decoder layers of LLaMA-2-7b are calculated and presented in Figure 1.
The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components. The distributions of their power spectrums are similar since they are projected from the same hidden states as shown in Equation 3. Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure. The observation of energy concentration suggests that we could maintain low-frequency components and filter out high-frequency components which could be redundant. Head-wise analysis of the power spectrum is provided in Appendix A.
4.2 KV Compression in the Frequency Domain
We conduct DCT along the sequence dimension to transfer the KV cache to the frequency domain:
${\bm{Y}}_{0:N-1}^{K}=\text{DCT}(\displaystyle{\bm{K}}_{0:N-1}),\quad\displaystyle{\bm{Y}}_{0:N-1}^{V}=\text{DCT}(\displaystyle{\bm{V}}_{0:N-1}).$ (6)
As observed in Figure 1, since the lower-frequency components are of higher magnitude and carry more information, we will retain them and remove higher-frequency components for compression. Given the retaining ratio $\gamma$, the retaining size is $L=\gamma\cdot N$. $N-L$ high-frequency components are filtered out to reduce redundancy:
$\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}=\displaystyle{\bm{Y}}_{0:N-1}^{K}[0:L-1],\quad\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}=\displaystyle{\bm{Y}}_{0:N-1}^{V}[0:L-1].$ (7)
Then we conduct IDCT along the frequency dimension to convert the compressed components back to the time dimension. It should be noted that the signals are normalized by the square root of the component number as shown in the formulas of DCT and IDCT. While DCT works on $N$ time-domain signals, only $L$ frequency-domain signals are processed by IDCT. The compressed states are amplified by $\sqrt{\frac{N}{L}}$ during the tranforms. Therefore, the compressed signals should be rescaled with the factor of $\sqrt{\frac{L}{N}}$ to restore the original amplitude:
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}),\quad\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}).$ (8)
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV cache of size $L$ in the time domain. The superscript “0:$N-1$” means that $\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV of the cached $N$ tokens. The subscript “0:$L-1$” means the retaining size is $L$. The incoming token $x_{N}$ will attend to the compressed KV cache:
$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(N,L)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{N}[\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{k}}_{N}]^{T}}{\sqrt{d}}\right)\cdot[\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{v}}_{N}].$ (9)
Figure 2: The illustration of our FreqKV. The KV cache will be compressed in an iterative manner as the cache reaches the context window size. Sink tokens remain uncompressed throughout the process. (a) The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache. (b) When the cache is filled again, the compressed tokens and incoming tokens will be compressed together. The compression is performed iteratively to extend the context window.
4.3 Context Extension with the Compressed KV
When extending the context window, the memory requirement of KV cache increases linearly and the computation cost of the original full attention grows quadratically with the length. To limit the size of KV cache for each query token to attend below the context window size, we will compress KV cache in the frequency domain iteratively as the cache is filled. We illustrate our FreqKV in Figure 2.
For tokens within the context window, the standard attention will be conducted. For tokens out of the window, we will compress the cached KV states in the frequency domain. Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache. We discard the original KV states of the prefilling tokens and maintain the compressed KV states. They will be concatenated with the KV states of the incoming tokens and compressed together when the cache is filled again. Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens. With the iterative FreqKV, the KV cache size is not fixed and is reduced below the context window size during decoding. Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible. For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens [26, 12]. Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
During training and the prefilling stage of inference, the whole sentence is tokenized and fed into the model. The attention is computed chunk-wise interleaved with the compression operation. After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed. As shown in Figure 2 (b), the incoming tokens in each chunk can not attend to the subsequent tokens. The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself. The calculation of attention in Equation 9 can be reformulated as follows:
$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot\quad[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$.
It should be noted that key states in LLMs are often equipped with position embeddings like RoPE [21]. In FreqKV, key states are compressed and cached before applying RoPE. The cached key states are encoded with RoPE when conducting self-attention. They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation. Compressing key states with RoPE is explored in Appendix B.
5 Experiments
5.1 Implementation
We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b [22] and LLaMA-3-8b [1]. We maintain $S=4$ sink tokens uncompressed. The retaining ratio $\gamma$ in compression is set to 0.5. The preset context window size of LLaMA-2 is 4096, which is also the maximum KV cache size $N$. Therefore, the retaining size during each compression is $L=\gamma\cdot(N-S)=2046$. As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states. The context size of LLaMA-3 is 8192. It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
Minimal training is introduced to adapt models to this frequency-domain compression method. For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset, extending the context window from 4K to 8K, 16K, and 32K. Perplexity (PPL) evaluation is conducted on PG-19 [20] and Proof-pile [2]. For long context understanding, the instruction following dataset LongAlpaca [7] is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct. The context window is extended from 4K to 8K, and from 8K to 16K respectively. Models are evaluated on LongBench [3] and Needle-in-a-Haystack [10]. Details of training settings are summarized in Appendix C.
Table 1: Perplexity evaluation on the test sets of PG-19 and Proof-pile. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from [7].
Training Length Method Inference Cache Evaluation Context Length
2048 4096 8192 16384 32768
PG-19
8192 Full FT Full 7.55 7.21 6.98 - -
LongLoRA Full 7.70 7.35 7.14 - -
LoCoCo* Compressed 8.15 8.08 7.27 - -
FreqKV Compressed 7.45 7.12 7.04 7.02 7.02
16384 LongLoRA Full 7.65 7.28 7.02 6.86 -
FreqKV Compressed 7.46 7.13 7.03 6.99 6.98
32768 LongLoRA Full 8.29 7.83 7.54 7.35 7.22
FreqKV Compressed 7.47 7.14 7.04 7.00 6.98
Proof-pile
8192 Full FT Full 3.14 2.85 2.66 - -
LongLoRA Full 3.20 2.91 2.72 - -
LoCoCo* Compressed 3.40 3.20 2.88 - -
FreqKV Compressed 3.15 2.88 2.79 2.76 2.75
16384 LongLoRA Full 3.17 2.87 2.66 2.51 -
FreqKV Compressed 3.15 2.88 2.78 2.74 2.73
32768 LongLoRA Full 3.35 3.01 2.78 2.61 2.50
FreqKV Compressed 3.16 2.89 2.78 2.74 2.72
5.2 Long Context Language Modeling
We use FreqKV to train LLaMA-2-base-7b on RedPajama with lengths of 8192, 16384, and 32768. Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256. We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA [7], and LoCoCo [5]. While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
PPL scores on different evaluation context lengths are reported in Table 1. While both of LongLoRA and LoCoCo sacrifice performance within the original context length (2K and 4K), FreqKV almost has no performance degradation. Even with the training length of 8K, FreqKV performs comparably to LongLoRA with the training length of 32K. It also outperforms LoCoCo on the extended context length as well as the shorter lengths.
It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths. Proof-pile consists of a diverse set of mathematical texts, including both informal and formal sources, making it particularly challenging for models using compressed caches to predict the next token. In contrast, LongLoRA retains the full KV cache, which naturally benefits next-token prediction in such structured domains. However, the performance of FreqKV does not deteriorate as the context length increases. Our compressed cache method even surpasses Full FT and LongLoRA on PG-19.
Furthermore, we evaluate the performance of FreqKV at longer lengths as shown in Figure 5. With the training length of 8K, it effectively extends the context window to 64K, which is 16 times larger than the original size.
5.3 Long Context Understanding
To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K. Models are evaluated on the long context understanding benchmark LongBench [3] and Needle-in-a-Haystack [10].
LongBench.
Scores on the 5 categories of LongBench are reported in Table 2. FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks. We compare our method with different KV compression strategies, including LM-Infinite [12], LongHeads [18], SnapKV [17] and PyramidKV [6]. Most baselines use LLaMA-2-chat-7b as their backbone. The retaining ratio is set to 0.5. FreqKV achieves SOTA (state-of-the-art) on 10 long context understanding tasks for LLaMA-2 and 8 tasks for LLaMA-3.
Needle-in-a-Haystack.
The Needle-in-a-Haystack results are illustrated in Figure 3. FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3. In contrast, PyramidKV [6] fails beyond the original window boundary. This is because it merely compresses the KV cache at the first decoding step. Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights. Consequently, those unseen positions lead to the performance collapse.
Table 2: Scores of different KV compression methods on LongBench. The superscript “*” means that we reproduce SnapKV and PyramidKV following their official codes for evaluation. The results of LM-Infinite and LongHeads are reported from [18].
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code
NtrvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8
+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55
+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22
+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14
+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17
+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65
LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84
+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33
+FreqKV 22.18 42.43 44.19 46.63 40.24 19.57 30.48 23.3 28.53 69.5 88.32 41.73 60.41 51.0
(a) Results of PyramidKV on LLaMA-2. The average accuracy is 46.7%.
(b) Results of Dropping on LLaMA-2. The average accuracy is 75.5%.
(c) Results of FreqKV on LLaMA-2. The average accuracy is 84.2%.
(d) Results of FreqKV on LLaMA-3. The average accuracy is 90.7%.
Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length (“haystack”) ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” (a short sentence) within the document.
To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b. Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM [26]. It shares the same training settings as FreqKV. While Dropping suffers from the permanent loss of earlier tokens, FreqKV could retrieve information from a specific sentence (the “needle”) hidden within a document (the “haystack”) and achieve a better performance.
6 Analysis
Figure 4: Perplexity evaluation on the validation set of PG-19 with different retaining ratios.
Figure 5: Decoding time with the full cache and FreqKV on the increasing sequence length.
6.1 Retaining Ratio
We train LLaMA-2-7b on RedPajama [8] with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
The performance of FreqKV with different retaining ratios is presented in Figure 5. FreqKV performs better with more frequency components retained. Employing an iterative compression manner, less compressions are performed on the more recent tokens by FreqKV, which benefits long context language modeling. It consistently achieves low perplexity as the evaluation context length grows from 2K to 64K, far exceeding the training length.
6.2 Decoding Latency
Table 3: The decoding latency (seconds) of FreqKV and Dropping on LLaMA-2-7b with the increasing sequence length. The difference between FreqKV and Dropping shows the compression latency introduced by FreqKV.
Models 4K 8K 16K 24K 32K
Compression Times 0 3 7 11 15
Dropping 123.78$\pm$1.46 253.32$\pm$5.76 506.60$\pm$5.11 760.46$\pm$7.15 1022.50$\pm$18.83
FreqKV 124.34$\pm$1.80 252.75$\pm$3.55 507.61$\pm$5.44 764.82$\pm$9.72 1029.00$\pm$19.97
FreqKV performs compressions on-the-fly during decoding. We compare the decoding time required by LLaMA-2-7b with the full cache and our FreqKV when the sequence length increases. As shown in Figure 5, the decoding time starts to diverge at the length of 4K. While the full cache utilization leads to a quadratic growth in decoding time, the decoding time of FreqKV increases approximately linearly with a negligible time spent on compression, showcasing its efficiency.
Furthermore, we measure the decoding latency of FreqKV and Dropping with the increasing length. Dropping removes the earliest tokens instead of compression in the frequency domain. The difference in the decoding time between the two methods shows the latency of compression. The sink size is set to 4 and the retaining ratio is 0.5. Table 3 reports the average decoding time (seconds) over five runs, along with standard deviations. It can be observed that the additional latency (6.5s) introduced by FreqKV accounts for only 0.64% of the total decoding time at 32K tokens. Further analysis for the compression overhead is provided in Appendix D.
7 Conclusion
In this paper, we introduce FreqKV to compress KV states iteratively in the frequency domain for LLMs. We exploit the energy concentration of KV states in the frequency domain within the decoder layers. Specifically, we filter out the high-frequency components that are of low magnitude and retain the low-frequency components for compression. The KV cache is compressed in the frequency domain without introducing additional compression modules. Iteratively compressing the KV cache, FreqKV could extend the context window efficiently for LLMs. With minimal training of low-rank adaption, LLMs learn to leverage the compressed KV cache. Through extensive experiments and analysis on long context modeling and understanding, FreqKV demonstrates its efficiency and effectiveness in context extension.
Appendix A Head-wise Analysis of Power Spectrums
We explore the power spectrum distribution of key states and value states in different attention heads of LLaMA-2-7b as shown in Figure 6. Although values of power spectrums vary in different heads, their distribution exhibits similar patterns. They have a consistent tendency to aggregate energy in the low-frequency components along the decoding procedure. It could be promising to study specific differences and associations in different heads or other modules.
(a) The average power spectrums of key states.
(b) The average power spectrums of value states.
Figure 6: The average power spectrums of key states and value states in different heads of Llama-2-7b.
Appendix B Compressing Key States with RoPE
Figure 7: The average power spectrums of key states with RoPE.
We further investigate to compress key states after applying RoPE. Power spectrums of key states with RoPE are provided in Figure 7. Components of certain frequencies are filtered out by RoPE.
We train LLaMA-2-7b on RedPajama with the variant method that key states are compressed after applying RoPE. The training length is 8K. Models are evaluated on the test set of PG-19 and Proof-pile. The results in Table 4 demonstrate that the variant suffers from performance degradation, which is probably due to the usage of position embeddings out of the original window. In contrast, FreqKV compresses and caches key states before applying RoPE. Therefore, they are allocated with position indices within the original window rather than the lengthy sequence, which results in the context window extension with no need of using out-of-bound position embeddings.
Table 4: Perplexity evaluation on the test sets of PG-19 and Proof-pile. “Variant” stands for FreqKV with key states compressed after applying RoPE.
Dataset Method Evaluation Context Length
2048 4096 8192
PG-19 Full FT 7.55 7.21 6.98
FreqKV 7.45 7.12 7.04
Variant 7.53 7.19 7.13
Proof-pile Full FT 3.14 2.85 2.66
FreqKV 3.15 2.88 2.79
Variant 3.16 2.88 2.80
Appendix C Training settings
For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset for 1000 steps. For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca [7] for 5 epochs, which consists of 6.28K long-context QA samples.
The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64. The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps. The rank used in the LoRA [15] fine-tuning is set to 8. Following LongLoRA [7], the embedding and normalization layers are learnable during training. All the experiments are conducted on ADA6000 and H100 GPUs. Moreover, we equip our method with FlashAttention-2 [9] for further acceleration and memory saving.
Appendix D Compression Overhead
We conduct further studies on the compression overhead of FreqKV. As introduced in Section 4.3, chunk-wise attention is performed for the prefilling tokens. The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens. Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention [4].
We use torchprofile [^1] https://github.com/zhijian-liu/torchprofile to count the number of Floating Point Operations (FLOPs) with input sequences of different lengths for LLaMA-2-7b. We calculate FLOPs for the model with the original attention which leverages full KV states, and with FreqKV with the retaining ratio of 0.5. To quantify the compression overhead, we have measured FLOPs of the Dropping, which retains the most recent tokens instead of compressing in the frequency domain. It shares the same sink size and retaining size as FreqKV. The difference in FLOPs between the two methods shows the overhead of compression.
The statistics are given in Table 5. The compression times of FreqKV with different context lengths are also reported in the table. It shows that the computation overhead of our compression process grows less than 0.5% even with a length of 32K, which could be negligible. FreqKV reduces more FLOPs as the input length grows from 4K to 32K compared to Full KV. This is because the compression is performed every $N-L-S$ tokens with the complexity of $O(N\text{log}N)$, which is negligible compared to the quadratic self-attention.
Table 5: FLOPs (TFLOPs) with input sequences of different lengths. Experiments are conducted on the ADA6000 GPU of 48GB, where full KV of 12K tokens with float16 will cause an OOM (Out-of-Memory) issue. The difference between FreqKV and Dropping shows the computation overhead of compression.
Models 4K 8K 12K 16K 32K
Full KV 62.93 143.46 OOM OOM OOM
Dropping 62.93 125.86 188.79 251.72 503.43
FreqKV 62.93 125.90 188.85 251.81 503.63
Compression Overhead 0 (0%) 0.039 (0.031%) 0.064 (0.034%) 0.090 (0.036%) 0.193 (0.038%)
Compression Times 0 3 5 7 15

### v2

@@
 Abstract
-Extending the context window in large language models (LLMs) is essential for applications involving long-form content generation.
-However, the linear increase in key-value (KV) cache memory requirements and the quadratic complexity of self-attention with respect to sequence length present significant challenges during fine-tuning and inference.
-Existing methods suffer from performance degradation when extending to longer contexts.
-In this work, we introduce a novel context extension method that optimizes both fine-tuning and inference efficiency.
-Our method exploits a key observation: in the frequency domain, the energy distribution of the KV cache is primarily concentrated in low-frequency components.
-By filtering out the high-frequency components, the KV cache can be effectively compressed with minimal information loss.
-Building on this insight, we propose an efficient compression technique, FreqKV, that iteratively compresses the increasing KV cache to a fixed size in the frequency domain, applicable to both fine-tuning and inference.
-FreqKV introduces no additional parameters or architectural modifications.
-With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window efficiently.
-Experiments on various long context language modeling and understanding tasks demonstrate the efficiency and efficacy of the proposed method.
+Frequency-domain compression has proven effective in reducing redundancies for spatial signals.
+In this work, we propose FreqKV, a novel frequency domain key-value (KV) compression technique that enables efficient context window extension for decoder-only large language models (LLMs).
+Our approach is motivated by a key observation that, in the frequency domain, the energy distribution of the KV cache is predominantly concentrated in low-frequency components.
+By discarding high-frequency components, we achieve efficient compression of the KV cache with minimal information loss.
+FreqKV iteratively compresses the increasing KV cache to a fixed size in the frequency domain, allowing models to process lengthy contexts efficiently.
+Introducing no additional parameters or architectural modifications, FreqKV is applicable to both fine-tuning and inference.
+With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window.
+Experiments on a range of long context language modeling and understanding tasks demonstrate the efficiency and effectiveness of the proposed method.
 1 Introduction
 Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process.
-However, it is inevitable for LLMs to process sequences that exceed the preset context size.
-LLMs struggle to maintain their performance when generalized to longer contexts.
-Additionally, the computation cost of the self-attention mechanism [23] grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
+However, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size.
+Additionally, the computational cost of the self-attention mechanism (Vaswani et al.,, 2023) grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
+Recent studies propose to fine-tune LLMs to longer contexts to extend the context window.
+LongLoRA (Chen et al.,, 2024) trains LLMs using shifted sparse attention.
+Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence.
+Concurrently, LoCoCo (Cai et al., 024a, ) and Activation Beacon (Zhang et al., 2024a, ) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
+However, the increasing memory requirement for additional parameters is inevitable.
 For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference.
-They evict [26, 17] or merge [29, 24] KV states of less important tokens following certain rules.
+They evict (Xiao et al.,, 2024; Li et al.,, 2024; Jo et al.,, 2025) or merge (Zhang et al., 2024b, ; Wan et al.,, 2024) KV states of less important tokens following certain rules.
 They use attention scores to measure the importance and approximate the original full attention.
-However, while these methods provide an approximation of full computation on existing tokens through different strategies, they can not fully prevent performance degradation when decoding future tokens.
-Recent studies propose to fine-tune LLMs to longer contexts to extend the context window.
-LongLoRA [7] trains LLMs using shifted sparse attention.
-Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence.
-Concurrently, LoCoCo [5] and Activation Beacon [28] introduce additional modules to compress KV states.
-They incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
-In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) [27].
-Moreover, Fourier Transformer [13] discards the high-frequency parts of the contexts and downsample the hidden states in the encoders.
+However, while these methods provide an approximation of full computation on existing tokens, the information associated with the evicted tokens will be lost.
+They can not fully prevent performance degradation when decoding future tokens.
+In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) (Xu et al.,, 2020).
+Moreover, Fourier Transformer (He et al.,, 2023) discards the high-frequency parts of the contexts and downsample the hidden states in the encoders.
 Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs.
-We transform key states and value states in LLaMA-2-7b [22] from the time domain to the frequency domain for power spectrum analysis.
-As shown in Figure 1, the energy distribution increasingly concentrates on low-frequency components as the computation process progresses along the decoder layers.
+We transform key states and value states in LLaMA-2-7b (Touvron et al.,, 2023) from the time domain to the frequency domain for power spectrum analysis.
+As shown in Figure 1, the energy distribution is increasingly concentrated on low-frequency components as the computation process progresses along the decoder layers.
 It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
-In this paper, we introduce FreqKV, an efficient context extension method that iteratively compresses key-value states in the frequency domain.
+In this paper, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain.
 Compression is triggered only when the KV cache reaches the predefined context window size.
-During each compression step, the low-frequency components of the KV states are preserved at a specified retaining ratio.
+During each compression step, KV states are converted to the frequency domain, where low-frequency components are retained before being transformed back to the time domain.
 Subsequent tokens are appended to the compressed cache until it is filled again.
-This ensures that the cached KV states that each query token can attend to is limited below the context window size.
-To reduce memory and computational costs, the compressed cache will be further compressed together with the incoming tokens.
-This iterative compression mechanism leads to an increased compression level of the earlier contexts as the sequence length grows.
+This ensures that the cached KV states that each query token can attend to are limited below the context window size, reducing memory and computational costs.
+The compressed cache undergoes further compression alongside incoming tokens.
+This iterative process progressively increases the compression ratio of earlier contexts as the sequence lengthens, while recent tokens remain less compressed, which is beneficial for autoregressive models to predicet next tokens.
 Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts.
 FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling.
-Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long-context understanding.
+Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long context understanding.
 2 Related Work
+Context Extension for LLMs.
+Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs.
+LongLoRA (Chen et al.,, 2024) employs shifted sparse attention during the parameter-efficient fine-tuning.
+However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training.
+Other techniques, such as LoCoCo (Cai et al., 024a, ), integrate convolutional operations into LLMs for compressing long contexts.
+They fine-tune the compression modules together with LLMs.
+Landmark attention (Mohtashami and Jaggi,, 2023) uses landmark tokens to retrieve previous input blocks.
+Similarly, Activation Beacon (Zhang et al., 2024a, ) introduces a special token to represent the previous context for compression.
+However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models.
+In contrast, our proposed method achieves context extension without introducing any additional parameters.
 KV Compression for LLMs.
 To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model.
-One common approach is selective token eviction [26, 17, 6], where less significant tokens are discarded.
+One common approach is selective token eviction (Xiao et al.,, 2024; Li et al.,, 2024; Cai.
+et al., 024b, ; Jo et al.,, 2025), where less significant tokens are discarded.
 Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens.
-To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts [29, 24, 25].
+To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts (Zhang et al., 2024b, ; Wan et al.,, 2024; Wang et al.,, 2024).
 However, these inference methods often sacrifice performance for efficiency.
-Context Extension for LLMs.
-Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs.
-LongLoRA [7] employs shifted sparse attention during the parameter-efficient fine-tuning.
-However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training.
-Other techniques, such as LoCoCo [5], integrate convolutional operations into LLMs for compressing long contexts.
-They fine-tune the compression modules together with LLMs.
-Landmark attention [19] uses landmark tokens to retrieve previous input blocks.
-Similarly, Activation Beacon [28] introduces a special token to represent the previous context for compression.
-However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models.
-In contrast, our proposed method achieves context extension without introducing any additional parameters.
 Learning in the Frequency Domain.
-Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs [11].
-It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels [27].
+Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al.,, 2018).
+It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al.,, 2020).
 These works have inspired efforts to process natural language.
-FNet [16] enhances the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens.
-Additionally, Fourier Transformer [13] eliminates redundancies in the context through frequency domain processing within encoder architectures.
-However, because of the auto-regressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs.
+FNet (Lee-Thorp et al.,, 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens.
+Additionally, Fourier Transformer (He et al.,, 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
+However, because of the autoregressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs.
 To the best of our knowledge, FreqKV is the first work that explores compressing key-value states in the frequency domain for decoder-only LLMs.
@@
 The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain.
-Several variants of the DCT exist, with DCT-II being the most common.
+Several variants of DCT exist, with DCT-II being the most common.
 For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
@@
 The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components.
-The components of higher energy in the frequency domain indicate they are more informative [13].
+The components of higher energy in the frequency domain indicate they are more informative (He et al.,, 2023).
 3.2 Self-Attention
-For the incoming token $x_{N}$, the prefilled $N$ tokens $X_{0:N-1}$ are utilized as the cache during decoding.
-Denote the hidden states of the $N+1$ tokens input to a specific layer of LLMs as ${\bm{H}}_{0:N}=[\displaystyle{\bm{h}}_{0},\cdots,\displaystyle{\bm{h}}_{N}]$.
-The query, key and value states of ${\bm{x}}_{N}$ are computed as follows::
-${\bm{q}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{Q},\quad\displaystyle{\bm{k}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{v}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{V},$ (3)
-where ${\bm{W}}^{Q},\displaystyle{\bm{W}}^{K},\displaystyle{\bm{W}}^{V}$ are the projection matrices for the query, key and value states, respectively.
+For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}$ are utilized as the cache during decoding.
+Denote the cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ as ${\bm{K}}_{0:N-1}$ and ${\bm{V}}_{0:N-1}$.
 For simplicity, indices corresponding to layers and heads have been omitted.
-The cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ are:
-${\bm{K}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{V}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{V}.$ (4)
 When calculating attention scores, the incoming token $x_{N}$ attends to all cached KV states as well as to itself:
 $\mathcal{\displaystyle{\bm{A}}}(N)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{N}[\displaystyle{\bm{K}}_{0:N-1}\oplus\displaystyle{\bm{k}}_{N}]^{T}}{\sqrt{d}}\right)\cdot[\displaystyle{\bm{V}}_{0:N-1}\oplus\displaystyle{\bm{v}}_{N}],$ (3)
-where $d$ is the hidden dimension. $\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$
+where $d$ is the hidden dimension.
+${\bm{q}}_{N},\displaystyle{\bm{k}}_{N}$ and ${\bm{v}}_{N}$ are the query, key and value states of $x_{N}$ respectively.
+$\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$.
 4 FreqKV
+In this section, we present power spectrum analysis of KV states for LLMs.
+Building on this observation, FreqKV compresses KV states in the frequency domain and extends the context window via iterative compression.
 (a) The average power spectrums of key states.
@@
 Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b.
-1000 documents are sampled from CNN/Daily Mail [14].
+1000 documents are sampled from CNN/Daily Mail (Hermann et al.,, 2015).
 We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
@@
 The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components.
-The distributions of their power spectrums are similar since they are projected from the same hidden states as shown in Equation 3.
+The distributions of their power spectrums are similar since they are projected from the same hidden states.
 Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure.
@@
 Sink tokens remain uncompressed throughout the process.
-(a) The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache.
-(b) When the cache is filled again, the compressed tokens and incoming tokens will be compressed together.
+The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache.
+When the cache is filled again, the compressed tokens and incoming tokens will be compressed together.
 The compression is performed iteratively to extend the context window.
-4.3 Context Extension with the Compressed KV
+4.3 Context Extension via Iterative Compression
 When extending the context window, the memory requirement of KV cache increases linearly and the computation cost of the original full attention grows quadratically with the length.
@@
 For tokens out of the window, we will compress the cached KV states in the frequency domain.
+During each compression step, KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain.
 Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache.
-We discard the original KV states of the prefilling tokens and maintain the compressed KV states.
-They will be concatenated with the KV states of the incoming tokens and compressed together when the cache is filled again.
+The compressed cache will be concatenated with KV states of the incoming tokens and compressed together when the cache is filled again.
 Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens.
-With the iterative FreqKV, the KV cache size is not fixed and is reduced below the context window size during decoding.
 Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible.
 For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
-Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens [26, 12].
+Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al.,, 2024; Han et al.,, 2024).
 Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
@@
 After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed.
-As shown in Figure 2 (b), the incoming tokens in each chunk can not attend to the subsequent tokens.
+As shown in Figure 2, the incoming tokens in each chunk can not attend to the subsequent tokens.
 The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself.
@@
 $\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot\quad[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$.
-It should be noted that key states in LLMs are often equipped with position embeddings like RoPE [21].
+It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al.,, 2023).
 In FreqKV, key states are compressed and cached before applying RoPE.
@@
 5.1 Implementation
-We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b [22] and LLaMA-3-8b [1].
+We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b (Touvron et al.,, 2023) and LLaMA-3-8b (AI@Meta,, 2024).
 We maintain $S=4$ sink tokens uncompressed.
@@
 As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states.
-The context size of LLaMA-3 is 8192.
+As for LLaMA-3, its context size is 8192.
 It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
 Minimal training is introduced to adapt models to this frequency-domain compression method.
-For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset, extending the context window from 4K to 8K, 16K, and 32K.
-Perplexity (PPL) evaluation is conducted on PG-19 [20] and Proof-pile [2].
-For long context understanding, the instruction following dataset LongAlpaca [7] is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct.
+For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer,, 2023), extending the context window from 4K to 8K, 16K, and 32K.
+Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al.,, 2019) and Proof-pile (Azerbayev et al.,, 2022).
+For long context understanding, the instruction following dataset LongAlpaca (Chen et al.,, 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct.
 The context window is extended from 4K to 8K, and from 8K to 16K respectively.
-Models are evaluated on LongBench [3] and Needle-in-a-Haystack [10].
+Models are evaluated on LongBench (Bai et al.,, 2024) and Needle-in-a-Haystack (gkamradt,, 2023).
 Details of training settings are summarized in Appendix C.
@@
 The superscript “*” means that we reproduce LoCoCo following their official code for evaluation.
-The results of full fine-tuning and LongLoRA are reported from [7].
+The results of full fine-tuning and LongLoRA are reported from (Chen et al.,, 2024).
 Training Length Method Inference Cache Evaluation Context Length
@@
 Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256.
-We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA [7], and LoCoCo [5].
+We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA (Chen et al.,, 2024), and LoCoCo (Cai et al., 024a, ).
 While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
@@
 It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths.
-Proof-pile consists of a diverse set of mathematical texts, including both informal and formal sources, making it particularly challenging for models using compressed caches to predict the next token.
-In contrast, LongLoRA retains the full KV cache, which naturally benefits next-token prediction in such structured domains.
+Proof-pile consists of a diverse set of mathematical texts, making it particularly challenging for models using compressed cache to predict the next token.
+In contrast, LongLoRA retains full KV cache during inference, which naturally benefits next-token prediction in such structured domains.
 However, the performance of FreqKV does not deteriorate as the context length increases.
@@
 To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K.
-Models are evaluated on the long context understanding benchmark LongBench [3] and Needle-in-a-Haystack [10].
+Models are evaluated on the long context understanding benchmark LongBench and Needle-in-a-Haystack.
 LongBench.
@@
 FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks.
-We compare our method with different KV compression strategies, including LM-Infinite [12], LongHeads [18], SnapKV [17] and PyramidKV [6].
+We compare our method with different KV compression strategies, including LM-Infinite (Han et al.,, 2024), LongHeads (Lu et al.,, 2024), FastKV (Jo et al.,, 2025), SnapKV (Li et al.,, 2024) and PyramidKV (Cai.
+et al., 024b, ).
 Most baselines use LLaMA-2-chat-7b as their backbone.
 The retaining ratio is set to 0.5.
-FreqKV achieves SOTA (state-of-the-art) on 10 long context understanding tasks for LLaMA-2 and 8 tasks for LLaMA-3.
+FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score.
 Needle-in-a-Haystack.
@@
 FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3.
-In contrast, PyramidKV [6] fails beyond the original window boundary.
-This is because it merely compresses the KV cache at the first decoding step.
+In contrast, PyramidKV fails beyond the original window boundary.
+This is because it compresses the long prompt after the first decoding step.
 Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights.
@@
 Table 2: Scores of different KV compression methods on LongBench.
-The superscript “*” means that we reproduce SnapKV and PyramidKV following their official codes for evaluation.
-The results of LM-Infinite and LongHeads are reported from [18].
-Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code
+The superscript “*” means that we reproduce FastKV, SnapKV and PyramidKV following their official codes for evaluation.
+The results of LM-Infinite and LongHeads are reported from Lu et al., (2024).
+Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
 NtrvQA
@@
 RB-P
-LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8
-+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55
-+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22
-+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14
-+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17
-+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65
-LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84
-+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33
-+FreqKV 22.18 42.43 44.19 46.63 40.24 19.57 30.48 23.3 28.53 69.5 88.32 41.73 60.41 51.0
+LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
++LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55 19.16
++LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22 32.13
++FastKV* 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
++SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
++PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
++FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
+LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
++PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
++FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
 (a) Results of PyramidKV on LLaMA-2. The average accuracy is 46.7%.
@@
 (d) Results of FreqKV on LLaMA-3. The average accuracy is 90.7%.
-Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length (“haystack”) ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” (a short sentence) within the document.
+Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” within the document.
 To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b.
-Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM [26].
+Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al.,, 2024).
 It shares the same training settings as FreqKV.
@@
 6.1 Retaining Ratio
-We train LLaMA-2-7b on RedPajama [8] with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
+We train LLaMA-2-7b on RedPajama (Computer,, 2023) with different retaining ratios.
+The training length is 8K.
+Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
 The performance of FreqKV with different retaining ratios is presented in Figure 5.
@@
 Appendix C Training settings
-For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset for 1000 steps.
-For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca [7] for 5 epochs, which consists of 6.28K long-context QA samples.
+For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer,, 2023) pre-training dataset for 1000 steps.
+For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al.,, 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
 The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64.
 The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps.
-The rank used in the LoRA [15] fine-tuning is set to 8.
-Following LongLoRA [7], the embedding and normalization layers are learnable during training.
+The rank used in the LoRA (Hu et al.,, 2021) fine-tuning is set to 8.
+Following LongLoRA (Chen et al.,, 2024), the embedding and normalization layers are learnable during training.
 All the experiments are conducted on ADA6000 and H100 GPUs.
-Moreover, we equip our method with FlashAttention-2 [9] for further acceleration and memory saving.
+Moreover, we equip our method with FlashAttention-2 (Dao,, 2023) for further acceleration and memory saving.
+The training time and memory requirements of FreqKV and Full FT on the RedPajama are reported in Table 5.
+For comparison, we set $Batch\_size\_per\_device$ to 1 and $Gradient\_accumulation\_steps$ to 8.
+8$\times$×ADA6000-48GB GPUs are used.
+Statistics show that FreqKV significantly reduces both training time and memory requirements, demonstrating its efficiency.
+Table 5: The training time and memory requirements of FreqKV and Full FT.
+Methods Training Length Time (hours) Memory (GB)
+Full FT 8K 28.88 44.04
+FreqKV 8K 17.07 24.58
+16K 38.57 34.00
+32K 89.35 45.37
 Appendix D Compression Overhead
@@
 The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens.
-Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention [4].
+Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention (Beltagy et al.,, 2020).
 We use torchprofile [^1] https://github.com/zhijian-liu/torchprofile to count the number of Floating Point Operations (FLOPs) with input sequences of different lengths for LLaMA-2-7b.

### v3

@@
-FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension
+FreqKV: Key-Value Compression in Frequency Domain for Context Window Extension
 Abstract
-Frequency-domain compression has proven effective in reducing redundancies for spatial signals.
-In this work, we propose FreqKV, a novel frequency domain key-value (KV) compression technique that enables efficient context window extension for decoder-only large language models (LLMs).
-Our approach is motivated by a key observation that, in the frequency domain, the energy distribution of the KV cache is predominantly concentrated in low-frequency components.
-By discarding high-frequency components, we achieve efficient compression of the KV cache with minimal information loss.
-FreqKV iteratively compresses the increasing KV cache to a fixed size in the frequency domain, allowing models to process lengthy contexts efficiently.
-Introducing no additional parameters or architectural modifications, FreqKV is applicable to both fine-tuning and inference.
-With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window.
-Experiments on a range of long context language modeling and understanding tasks demonstrate the efficiency and effectiveness of the proposed method.
+Existing key-value (KV) cache compression methods for large language models (LLMs) often rely on token eviction, which risks losing critical local information in both long prefilling and decoding scenarios.
+When extrapolating beyond the pretrained context length, their performance degrades sharply on long-context benchmarks.
+Motivated by the observation in the frequency domain that the context information is concentrated in the low-frequency components, we propose FreqKV, a parameter-free and architecture-agnostic approach.
+It iteratively compresses the increasing KV cache in the frequency domain, allowing models to process lengthy contexts efficiently.
+With minimal training at 8K length, FreqKV extends the context window of LLaMA-2-7B up to 256K tokens while maintaining stable perplexity.
+Extensive experiments across prefilling and decoding demonstrate that FreqKV enables robust context window extension and consistently outperforms existing KV cache compression methods on LLaMA-2 and LLaMA-3, highlighting its effectiveness for both understanding and generation in long contexts.
+Our code is available at https://github.com/LUMIA-Group/FreqKV.
 1 Introduction
-Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process.
-However, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size.
-Additionally, the computational cost of the self-attention mechanism (Vaswani et al.,, 2023) grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
-Recent studies propose to fine-tune LLMs to longer contexts to extend the context window.
-LongLoRA (Chen et al.,, 2024) trains LLMs using shifted sparse attention.
+Large language models (LLMs) have demonstrated remarkable performance in natural language understanding and generation.
+However, their inference capabilities are fundamentally constrained by a pre-defined context window size established during pre-training.
+Consequently, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size, significantly limiting their applicability in long-context scenarios.
+To alleviate this limitation, several position encoding methods, such as ALiBi (Press et al., 2022), PI (Chen et al., 2023), and LongRoPE (Ding et al., 2024) have been proposed to extend the context window.
+Relying on full self-attention, the quadratic computational cost renders them prohibitively expensive at scale.
+LongLoRA (Chen et al., 2024) trains LLMs using shifted sparse attention.
 Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence.
-Concurrently, LoCoCo (Cai et al., 024a, ) and Activation Beacon (Zhang et al., 2024a, ) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
+A popular strategy to improve efficiency is to compress the key-value (KV) cache during inference.
+Methods such as SnapKV (Li et al., 2024), PyramidKV (Cai.
+et al., 2024b), and FastKV (Jo et al., 2025) evict tokens deemed less important according to attention scores.
+However, these compression strategies cannot be extrapolated beyond the context window.
+The information associated with the evicted tokens will be lost, leading to performance degradation when decoding future tokens.
+Another line of work (Zhang et al., 2024; Wan et al., 2025) attempts to merge tokens rather than dropping them, preserving more information.
+Nevertheless, their performance remains suboptimal without fine-tuning.
+Concurrently, LoCoCo (Cai et al., 2024) and Activation Beacon (Zhang et al., 2025) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
 However, the increasing memory requirement for additional parameters is inevitable.
-For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference.
-They evict (Xiao et al.,, 2024; Li et al.,, 2024; Jo et al.,, 2025) or merge (Zhang et al., 2024b, ; Wan et al.,, 2024) KV states of less important tokens following certain rules.
-They use attention scores to measure the importance and approximate the original full attention.
-However, while these methods provide an approximation of full computation on existing tokens, the information associated with the evicted tokens will be lost.
-They can not fully prevent performance degradation when decoding future tokens.
-In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) (Xu et al.,, 2020).
-Moreover, Fourier Transformer (He et al.,, 2023) discards the high-frequency parts of the contexts and downsample the hidden states in the encoders.
-Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs.
-We transform key states and value states in LLaMA-2-7b (Touvron et al.,, 2023) from the time domain to the frequency domain for power spectrum analysis.
-As shown in Figure 1, the energy distribution is increasingly concentrated on low-frequency components as the computation process progresses along the decoder layers.
-It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
-In this paper, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain.
-Compression is triggered only when the KV cache reaches the predefined context window size.
-During each compression step, KV states are converted to the frequency domain, where low-frequency components are retained before being transformed back to the time domain.
-Subsequent tokens are appended to the compressed cache until it is filled again.
-This ensures that the cached KV states that each query token can attend to are limited below the context window size, reducing memory and computational costs.
-The compressed cache undergoes further compression alongside incoming tokens.
-This iterative process progressively increases the compression ratio of earlier contexts as the sequence lengthens, while recent tokens remain less compressed, which is beneficial for autoregressive models to predicet next tokens.
-Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts.
-FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling.
-Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long context understanding.
+In this work, we observe that the KV states of LLMs exhibit strong energy concentration in the low-frequency components of the frequency domain, suggesting that high-frequency parts are largely redundant.
+Building on this observation, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain.
+Unlike eviction methods, FreqKV preserves the information of all tokens by transforming KV states into the frequency domain and discarding high-frequency components that carry limited information.
+By compressing iteratively as the cache grows, FreqKV retains recent tokens with minimal loss while progressively compressing earlier ones.
+This parameter-free method requires no architectural modifications and can be applied to both the prefilling stage (long-context understanding) and the decoding stage (long-context language modeling and generation).
+Extensive experiments show that FreqKV enables robust context window extension.
+With minimal fine-tuning at 8K length, FreqKV maintains low perplexity as the evaluation context length scales to 256K.
+It achieves comparable performance to methods that employ full cache or additional compressors in long context language modeling.
+Experimental results across LongBench, RULER, and Needle-in-a-Haystack indicate that FreqKV surpasses recently studied KV compression methods in the prefilling stage.
+Furthermore, FreqKV also delivers substantial improvements in long-generation capability on LongGenBench, validating its effectiveness in decoding scenarios.
+These results establish frequency-domain compression as a viable approach for extending the context window.
 2 Related Work
-Context Extension for LLMs.
+KV Compression for LLMs.
+To alleviate memory and computation costs, a growing body of research explores compressing the KV cache when processing long contexts.
+One widely studied approach is selective token eviction, which can be traced back to attention sinks (Xiao et al., 2024a; b), suggesting that preserving a small set of initial tokens helps stabilize model performance.
+Furthermore, MInference (Jiang et al., 2024) propose to accelerate prefilling with unique patterns in long-context attention matrices.
+More recent methods, SnapKV (Li et al., 2024), InfiniPot (Kim et al., 2024), PyramidKV (Cai.
+et al., 2024b), GemFilter (Shi et al., 2024), and FastKV (Jo et al., 2025), use attention scores to measure the importance of tokens and discard less important ones.
+However, LLMs suffer from the permanent loss of the information associated with evicted tokens during decoding.
+To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts, such as CaM (Zhang et al., 2024), KVMerger (Wang et al., 2024), and D2O (Wan et al., 2025).
+Nevertheless, these inference methods often sacrifice performance for efficiency and cannot be extrapolated beyond the context window.
+Efficient Context Extension for LLMs.
 Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs.
-LongLoRA (Chen et al.,, 2024) employs shifted sparse attention during the parameter-efficient fine-tuning.
+LongLoRA (Chen et al., 2024) employs shifted sparse attention during the parameter-efficient fine-tuning.
 However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training.
-Other techniques, such as LoCoCo (Cai et al., 024a, ), integrate convolutional operations into LLMs for compressing long contexts.
+Other techniques, such as LoCoCo (Cai et al., 2024), integrate convolutional operations into LLMs for compressing long contexts.
 They fine-tune the compression modules together with LLMs.
-Landmark attention (Mohtashami and Jaggi,, 2023) uses landmark tokens to retrieve previous input blocks.
-Similarly, Activation Beacon (Zhang et al., 2024a, ) introduces a special token to represent the previous context for compression.
+Landmark attention (Mohtashami & Jaggi, 2023) uses landmark tokens to retrieve previous input blocks.
+Similarly, Activation Beacon (Zhang et al., 2025) introduces a special token to represent the previous context for compression.
 However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models.
 In contrast, our proposed method achieves context extension without introducing any additional parameters.
-KV Compression for LLMs.
-To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model.
-One common approach is selective token eviction (Xiao et al.,, 2024; Li et al.,, 2024; Cai.
-et al., 024b, ; Jo et al.,, 2025), where less significant tokens are discarded.
-Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens.
-To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts (Zhang et al., 2024b, ; Wan et al.,, 2024; Wang et al.,, 2024).
-However, these inference methods often sacrifice performance for efficiency.
 Learning in the Frequency Domain.
-Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al.,, 2018).
-It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al.,, 2020).
+Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al., 2018).
+It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al., 2020).
+Moreover, Frequency Principle (Xu et al., 2019; Luo et al., 2019) theoretically shows that low frequencies converge faster than high frequencies for deep neural networks.
 These works have inspired efforts to process natural language.
-FNet (Lee-Thorp et al.,, 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens.
-Additionally, Fourier Transformer (He et al.,, 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
+FNet (Lee-Thorp et al., 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens.
+Additionally, Fourier Transformer (He et al., 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
 However, because of the autoregressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs.
 To the best of our knowledge, FreqKV is the first work that explores compressing key-value states in the frequency domain for decoder-only LLMs.
-3 Preliminaries
-3.1 Discrete Cosine Transform
-The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain.
-Several variants of DCT exist, with DCT-II being the most common.
-For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
-$y_{t}=\alpha_{t}\sum_{n=0}^{N-1}x_{n}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right],$ $\alpha_{t}=\begin{cases}\sqrt{\frac{1}{N}}&\text{ if }t=0,\\ \sqrt{\frac{2}{N}}&\text{ otherwise }\end{cases}$ (1)
-where $t=0,1,\cdots,N-1$. $\alpha_{t}$ is the normalization factor.
-The time-domain signal ${\bm{X}}_{0:N-1}$ can be recovered by applying the inverse DCT (IDCT) on the frequency components ${\bm{Y}}_{0:N-1}$:
-$x_{n}=\sum_{t=0}^{N-1}\alpha_{t}\cdot y_{t}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right].$ (2)
-The frequency components are expressed as a combination of the original signals.
-The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$.
-The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components.
-The components of higher energy in the frequency domain indicate they are more informative (He et al.,, 2023).
-3.2 Self-Attention
-For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}$ are utilized as the cache during decoding.
+3 FreqKV
+In this section, we present power spectrum analysis of KV states for LLMs.
+Building on this observation, FreqKV compresses KV states in the frequency domain and extends the context window via iterative compression.
+(a) The average power spectrums of key states.
+(b) The average power spectrums of value states.
+Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b.
+1000 documents are sampled from CNN/Daily Mail (Hermann et al., 2015).
+We transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
+3.1 Preliminaries
+Energy Concentration in the Frequency Domain.
+We use DCT (Discrete Cosine Transform) to transform key states and value states from the time domain, which is the sequence dimension, to the frequency domain.
+The formulations of DCT and the inverse DCT (IDCT) can be referred to in Appendix A.
+The average power spectrums in different decoder layers of LLaMA-2-7B are calculated and presented in Figure 1.
+The figure reveals that the energy of key states and value states is increasingly concentrated in the low-frequency components.
+The distributions of their power spectrums are similar since they are projected from the same hidden states.
+While the initial embeddings of natural languages in the first layer exhibit no strong low-frequency bias, subsequent layers gradually shift energy toward low frequencies along the decoding procedure.
+This observation suggests that the low-frequency components capture most of the essential information, whereas high-frequency parts carry limited signal and can be discarded with minimal loss.
+Further analysis of frequency components is provided in Appendix B.
+Self-Attention with KV Cache.
+For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ are utilized as the cache during decoding.
 Denote the cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ as ${\bm{K}}_{0:N-1}$ and ${\bm{V}}_{0:N-1}$.
@@
 $\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$.
-4 FreqKV
-In this section, we present power spectrum analysis of KV states for LLMs.
-Building on this observation, FreqKV compresses KV states in the frequency domain and extends the context window via iterative compression.
-(a) The average power spectrums of key states.
-(b) The average power spectrums of value states.
-Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b.
-1000 documents are sampled from CNN/Daily Mail (Hermann et al.,, 2015).
-We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
-4.1 Energy Concentration in the Frequency Domain
-We transform key states and value states from the time domain, which is the sequence dimension, to the frequency domain.
-The average power spectrums in different decoder layers of LLaMA-2-7b are calculated and presented in Figure 1.
-The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components.
-The distributions of their power spectrums are similar since they are projected from the same hidden states.
-Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure.
-The observation of energy concentration suggests that we could maintain low-frequency components and filter out high-frequency components which could be redundant.
-Head-wise analysis of the power spectrum is provided in Appendix A.
-4.2 KV Compression in the Frequency Domain
-We conduct DCT along the sequence dimension to transfer the KV cache to the frequency domain:
-${\bm{Y}}_{0:N-1}^{K}=\text{DCT}(\displaystyle{\bm{K}}_{0:N-1}),\quad\displaystyle{\bm{Y}}_{0:N-1}^{V}=\text{DCT}(\displaystyle{\bm{V}}_{0:N-1}).$ (4)
-As observed in Figure 1, since the lower-frequency components are of higher magnitude and carry more information, we will retain them and remove higher-frequency components for compression.
-Given the retaining ratio $\gamma$, the retaining size is $L=\gamma\cdot N$.
-$N-L$ high-frequency components are filtered out to reduce redundancy:
-$\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}=\displaystyle{\bm{Y}}_{0:N-1}^{K}[0:L-1],\quad\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}=\displaystyle{\bm{Y}}_{0:N-1}^{V}[0:L-1].$ (5)
-Then we conduct IDCT along the frequency dimension to convert the compressed components back to the time dimension.
-It should be noted that the signals are normalized by the square root of the component number as shown in the formulas of DCT and IDCT.
-While DCT works on $N$ time-domain signals, only $L$ frequency-domain signals are processed by IDCT.
-The compressed states are amplified by $\sqrt{\frac{N}{L}}$ during the tranforms.
-Therefore, the compressed signals should be rescaled with the factor of $\sqrt{\frac{L}{N}}$ to restore the original amplitude:
-$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}),\quad\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}).$ (6)
-$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV cache of size $L$ in the time domain.
-The superscript “0:$N-1$” means that $\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV of the cached $N$ tokens.
-The subscript “0:$L-1$” means the retaining size is $L$.
-The incoming token $x_{N}$ will attend to the compressed KV cache:
-$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(N,L)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{N}[\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{k}}_{N}]^{T}}{\sqrt{d}}\right)\cdot[\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{v}}_{N}].$ (7)
-Figure 2: The illustration of our FreqKV.
-The KV cache will be compressed in an iterative manner as the cache reaches the context window size.
+Figure 2: Cosine Similarity between perturbed and original inputs for different frequency components of KV states.
+3.2 Effect of frequency components
+To empirically investigate what information different frequency bands encode, we have conducted a targeted perturbation experiment.
+We sample 100 CNN/DailyMail documents and introduce controlled perturbations by randomly replacing 1% of words with their synonyms.
+We then transform the KV states of LLaMA-2 into the frequency domain and compare the cosine similarity between perturbed and original inputs for both low-frequency (lowest 50%) and high-frequency (highest 50%) components.
+The results across different layers are presented in Figure 2:
+It shows that the output states are robust to the perturbation with low-frequency components retained while high-frequency ones are more sensitive, supporting the interpretation that low frequencies encode global information, while high frequencies capture local details.
+We further evaluate this distinction on the summarization track of the downstream benchmark LongBench, comparing the performance of llama-2-chat-7b with low and high frequencies retained in Table 1.
+Summarization requires capturing global semantic structure and long-range coherence.
+Accordingly, retaining low-frequency components yields substantially better performance.
+Table 1: Performance on summarization tasks with high/low-frequency components retained.
+Components Retained GovReport QMSum MultiNews
+high-frequency 14.21 16.54 15.55
+low-frequency 25.51 21.81 26.9
+Together, these results strengthen our motivation: low-frequency components primarily encode global semantic contexts and long-range dependencies, whereas high-frequency components encode local details, which explains why preserving low-frequency information is both more robust and more beneficial for long-context understanding.
+Figure 3: The overview of our FreqKV.
+(a) The illustration of the frequency-domain compression.
+(b) The KV cache will be compressed in an iterative manner to extend the context window.
 Sink tokens remain uncompressed throughout the process.
@@
 When the cache is filled again, the compressed tokens and incoming tokens will be compressed together.
-The compression is performed iteratively to extend the context window.
+3.3 KV Compression in the Frequency Domain
+To reduce redundancy in the key-value (KV) cache, we compress KV states in the frequency domain as shown in Figure 3a.
+Specifically, we conduct DCT along the sequence dimension to transfer the KV cache to the frequency domain:
+${\bm{Z}}_{0:N-1}^{K}=\text{DCT}(\displaystyle{\bm{K}}_{0:N-1}),\quad\displaystyle{\bm{Z}}_{0:N-1}^{V}=\text{DCT}(\displaystyle{\bm{V}}_{0:N-1}),$ (2)
+where ${\bm{Z}}_{0:N-1}^{K}$ and ${\bm{Z}}_{0:N-1}^{V}$ are the spectral representations of the KV cache.
+As illustrated in Figure 1, since the low-frequency components exhibit higher magnitudes and carry more information, we will retain them and remove the high-frequency parts for compression.
+Given a retaining ratio $\gamma$, the cache size is reduced to $L=\gamma\cdot N$ by filtering out NLN-L high-frequency components.
+Then we conduct IDCT along the frequency dimension to convert the compressed components back to the time dimension.
+It should be noted that the signals are normalized by the square root of the component number when applying DCT and IDCT (Appendix A).
+Therefore, the compressed signals should be rescaled with the factor of $\sqrt{\frac{L}{N}}$ to restore the original amplitude:
+$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\displaystyle{\bm{Z}}_{0:N-1}^{K}[0:L-1]),\quad\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\displaystyle{\bm{Z}}_{0:N-1}^{V}[0:L-1]),$ (3)
+$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV cache of size $L$ in the time domain.
+The superscript “0:N1N-1” indicates the original token range, while the subscript “0:L1L-1” means the retaining size.
+The incoming token $x_{N}$ will attend to the compressed KV cache as follows:
+$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(N,L)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{N}[\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{k}}_{N}]^{T}}{\sqrt{d}}\right)\cdot[\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}\oplus\displaystyle{\bm{v}}_{N}].$ (4)
 3.4 Context Extension via Iterative Compression
-When extending the context window, the memory requirement of KV cache increases linearly and the computation cost of the original full attention grows quadratically with the length.
-To limit the size of KV cache for each query token to attend below the context window size, we will compress KV cache in the frequency domain iteratively as the cache is filled.
-We illustrate our FreqKV in Figure 2.
+Extending the context window of LLMs is fundamentally constrained by memory and computation cost.
+To address this, FreqKV employs an iterative compression strategy in the frequency domain that constrains the effective cache size while enabling processing of arbitrarily long sequences.
+The overall pipeline is illustrated in Figure 3b.
 For tokens within the context window, the standard attention will be conducted.
-For tokens out of the window, we will compress the cached KV states in the frequency domain.
-During each compression step, KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain.
-Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache.
-The compressed cache will be concatenated with KV states of the incoming tokens and compressed together when the cache is filled again.
-Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens.
-Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible.
-For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
-Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al.,, 2024; Han et al.,, 2024).
+Once the cache reaches the maximum size, we will compress it in the frequency domain.
+Specifically, cached KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain.
+The incoming tokens are appended to the compressed cache, and the process repeats when the cache is filled again.
+As a result, earlier tokens experience more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens—a property well aligned with the autoregressive nature of LLMs.
+The compression is triggered only once every NLN-L tokens, with the complexity of $O(NlogN)$ for each compression.
+Therefore, the computational overhead of the compression could be negligible.
+Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al., 2024b; Han et al., 2024).
 Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
-During training and the prefilling stage of inference, the whole sentence is tokenized and fed into the model.
-The attention is computed chunk-wise interleaved with the compression operation.
-After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed.
-As shown in Figure 2, the incoming tokens in each chunk can not attend to the subsequent tokens.
-The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself.
-The calculation of attention in Equation 7 can be reformulated as follows:
-$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot\quad[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$.
-It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al.,, 2023).
+During training and prefilling, FreqKV will process the sentence chunk-wise, interleaving attention computation with the compression operation.
+After each compression, NLSN-L-S incoming tokens form a chunk that fills the cache, while $S$ sink tokens remain uncompressed, and $L$ compressed “tokens” represent earlier history.
+Consequently, the newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed tokens, MNM-N previous incoming tokens, and itself:
+$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$,
+where $\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}$and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}$ denote the compressed KV cache of size $L$ for tokens ${\bm{X}}_{S:N-1}=[x_{S},\dots,x_{N-1}]$.
+It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al., 2024).
 In FreqKV, key states are compressed and cached before applying RoPE.
-The cached key states are encoded with RoPE when conducting self-attention.
+The cached key states are then encoded with RoPE when conducting self-attention.
 They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation.
-Compressing key states with RoPE is explored in Appendix B.
+Further discussion of RoPE integration is provided in Appendix C.
 4 Experiments
-5.1 Implementation
-We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b (Touvron et al.,, 2023) and LLaMA-3-8b (AI@Meta,, 2024).
-We maintain $S=4$ sink tokens uncompressed.
-The retaining ratio $\gamma$ in compression is set to 0.5.
-The preset context window size of LLaMA-2 is 4096, which is also the maximum KV cache size $N$.
-Therefore, the retaining size during each compression is $L=\gamma\cdot(N-S)=2046$.
-As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states.
-As for LLaMA-3, its context size is 8192.
-It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
-Minimal training is introduced to adapt models to this frequency-domain compression method.
-For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer,, 2023), extending the context window from 4K to 8K, 16K, and 32K.
-Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al.,, 2019) and Proof-pile (Azerbayev et al.,, 2022).
-For long context understanding, the instruction following dataset LongAlpaca (Chen et al.,, 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct.
+4.1 Experimental Settings
+We employ FreqKV on LLaMA-2-7B/13B (Touvron et al., 2023) and LLaMA-3-8B (AI@Meta, 2024).
+Unless otherwise specified, the retaining ratio $\gamma$ is set to 0.5, and $S=4$ sink tokens are kept uncompressed.
+For LLaMA-2, the preset context window size is 4096, which also defines the maximum KV cache size $N$.
+Accordingly, during each compression step, the NS=4092N-S=4092 states since the $5$-th state in the cache will be compressed into $L=\gamma\cdot(N-S)=2046$ states.
+As for LLaMA-3, the native context window size is 8192.
+It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2 (Multi-Head Attention, MHA).
+Minimal training is introduced to adapt models to this frequency-domain compression method, following LongLoRA (Chen et al., 2024).
+Details of training settings are summarized in Appendix D.
+We assess FreqKV on long context language modeling and generation tasks for decoding performance, and long context understanding tasks for prefilling performance.
+For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer, 2023), extending the context window from 4K to 32K.
+Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al., 2019).
+For long context generation and understanding, the instruction following dataset LongAlpaca (Chen et al., 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct.
 The context window is extended from 4K to 8K, and from 8K to 16K respectively.
-Models are evaluated on LongBench (Bai et al.,, 2024) and Needle-in-a-Haystack (gkamradt,, 2023).
-Details of training settings are summarized in Appendix C.
-Table 1: Perplexity evaluation on the test sets of PG-19 and Proof-pile.
+Models are evaluated on LongGenBench (Liu et al., 2024), LongBench (Bai et al., 2024), RULER (Hsieh et al., 2024), and Needle-in-a-Haystack (gkamradt, 2023).
+Table 2: Perplexity evaluation on the test set of PG-19.
 The superscript “*” means that we reproduce LoCoCo following their official code for evaluation.
-The results of full fine-tuning and LongLoRA are reported from (Chen et al.,, 2024).
-Training Length Method Inference Cache Evaluation Context Length
+The results of full fine-tuning and LongLoRA are reported from LongLoRA (Chen et al., 2024).
+Full FT and LoCoCo encounter the OOM (Out-of-Memory) issue with the training length of 16K.
+Size Training Length Method Inference Cache Evaluation Context Length
 2048 4096 8192 16384 32768
-PG-19
-8192 Full FT Full 7.55 7.21 6.98 - -
+7B 8192 Full FT Full 7.55 7.21 6.98 - -
+LoCoCo* Compressed 8.15 8.08 7.27 - -
 LongLoRA Full 7.70 7.35 7.14 - -
-LoCoCo* Compressed 8.15 8.08 7.27 - -
 FreqKV Compressed 7.45 7.12 7.04 7.02 7.02
-16384 LongLoRA Full 7.65 7.28 7.02 6.86 -
+16384 Full FT Full OOM during Training
+LoCoCo Compressed OOM during Training
+LongLoRA Full 7.65 7.28 7.02 6.86 -
 FreqKV Compressed 7.46 7.13 7.03 6.99 6.98
@@
 FreqKV Compressed 7.47 7.14 7.04 7.00 6.98
-Proof-pile
-8192 Full FT Full 3.14 2.85 2.66 - -
-LongLoRA Full 3.20 2.91 2.72 - -
-LoCoCo* Compressed 3.40 3.20 2.88 - -
-FreqKV Compressed 3.15 2.88 2.79 2.76 2.75
-16384 LongLoRA Full 3.17 2.87 2.66 2.51 -
-FreqKV Compressed 3.15 2.88 2.78 2.74 2.73
-32768 LongLoRA Full 3.35 3.01 2.78 2.61 2.50
-FreqKV Compressed 3.16 2.89 2.78 2.74 2.72
+13B 8192 Full FT Full 6.95 6.60 6.43 - -
+LongLoRA Full 7.03 6.73 6.58 - -
+FreqKV Compressed 6.82 6.52 6.44 6.42 6.41
 4.2 Long Context Language Modeling
-We use FreqKV to train LLaMA-2-base-7b on RedPajama with lengths of 8192, 16384, and 32768.
-Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256.
-We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA (Chen et al.,, 2024), and LoCoCo (Cai et al., 024a, ).
-While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
+We use FreqKV to train LLaMA-2-7B and LLaMA-2-13B on RedPajama.
+Perplexity is measured on the test set of the book corpus dataset PG-19 with the evaluation sliding window of 256.
+We compare our method with other baselines, including full fine-tuning (Full FT), LongLoRA, and LoCoCo (Cai et al., 2024).
+While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo and FreqKV use compressed cache.
 PPL scores on different evaluation context lengths are reported in Table 2.
-While both of LongLoRA and LoCoCo sacrifice performance within the original context length (2K and 4K), FreqKV almost has no performance degradation.
-Even with the training length of 8K, FreqKV performs comparably to LongLoRA with the training length of 32K.
-It also outperforms LoCoCo on the extended context length as well as the shorter lengths.
-It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths.
-Proof-pile consists of a diverse set of mathematical texts, making it particularly challenging for models using compressed cache to predict the next token.
-In contrast, LongLoRA retains full KV cache during inference, which naturally benefits next-token prediction in such structured domains.
-However, the performance of FreqKV does not deteriorate as the context length increases.
-Our compressed cache method even surpasses Full FT and LongLoRA on PG-19.
-Furthermore, we evaluate the performance of FreqKV at longer lengths as shown in Figure 5.
-With the training length of 8K, it effectively extends the context window to 64K, which is 16 times larger than the original size.
-5.3 Long Context Understanding
-To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K.
-Models are evaluated on the long context understanding benchmark LongBench and Needle-in-a-Haystack.
-LongBench.
-Scores on the 5 categories of LongBench are reported in Table 2.
-FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks.
-We compare our method with different KV compression strategies, including LM-Infinite (Han et al.,, 2024), LongHeads (Lu et al.,, 2024), FastKV (Jo et al.,, 2025), SnapKV (Li et al.,, 2024) and PyramidKV (Cai.
-et al., 024b, ).
-Most baselines use LLaMA-2-chat-7b as their backbone.
-The retaining ratio is set to 0.5.
-FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score.
-Needle-in-a-Haystack.
-The Needle-in-a-Haystack results are illustrated in Figure 3.
-FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3.
-In contrast, PyramidKV fails beyond the original window boundary.
-This is because it compresses the long prompt after the first decoding step.
-Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights.
-Consequently, those unseen positions lead to the performance collapse.
+More Evaluation results on Proof-pile (Azerbayev et al., 2022) are presented in Appendix E.
+While both LongLoRA and LoCoCo sacrifice performance within the original context length (2K and 4K), FreqKV matches or even surpasses Full FT.
+With the training length of 8K, FreqKV performs comparably to LongLoRA trained at 32K and consistently outperforms LoCoCo in terms of extended context length as well as the shorter lengths.
+Additional ablation studies on retained frequency components, retaining ratio, and sink size are reported in Appendix F.
 Table 3: Scores of different KV compression methods on LongBench.
-The superscript “*” means that we reproduce FastKV, SnapKV and PyramidKV following their official codes for evaluation.
-The results of LM-Infinite and LongHeads are reported from Lu et al., (2024).
-Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
+The superscript “*” means that results of CaM and KVMerger are reported from KVMerger (Wang et al., 2024).
+We reproduce the other baselines following their official codes for evaluation.
+The second-best scores are also underlined in the table.
+Retaining Ratio Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
 NtrvQA
@@
 RB-P
-LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
-+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55 19.16
-+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22 32.13
-+FastKV* 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
-+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
-+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
-+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
-LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
-+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
-+FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
-(a) Results of PyramidKV on LLaMA-2. The average accuracy is 46.7%.
-(b) Results of Dropping on LLaMA-2. The average accuracy is 75.5%.
-(c) Results of FreqKV on LLaMA-2. The average accuracy is 84.2%.
-(d) Results of FreqKV on LLaMA-3. The average accuracy is 90.7%.
-Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” within the document.
-To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b.
-Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al.,, 2024).
-It shares the same training settings as FreqKV.
-While Dropping suffers from the permanent loss of earlier tokens, FreqKV could retrieve information from a specific sentence (the “needle”) hidden within a document (the “haystack”) and achieve a better performance.
+LLaMA-2-chat-7B
+100% Full Cache 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
+50% CaM* 17.08 20.00 33.98 - 30.69 - 24.46 - 24.66 63.5 82.17 - - - -
+KVMerger* 18.50 20.04 36.89 - 32.99 - 25.31 - 26.29 64.0 83.62 - - - -
+LaCache 16.27 16.55 31.55 32.62 26.22 7.72 22.84 20.34 25.16 65.0 83.28 38.46 47.97 43.41 34.10
+D2O 20.32 19.16 33.09 30.62 26.76 9.65 23.72 20.54 25.42 63.0 84.23 41.22 58.61 52.77 36.37
+GemFilter 15.22 17.76 32.92 27.29 26.76 8.66 16.17 16.29 15.95 57.5 84.8 31.96 46.54 45.14 31.64
+FastKV 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
+SnapKV 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
+ThinK 15.57 21.76 37.98 30.21 31.55 8.06 26.24 20.98 26.1 65.0 80.08 40.97 60.2 54.26 37.07
+PyramidKV 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
+1% LaCache 4.37 12.72 6.34 14.77 14.91 2.23 11.18 7.39 9.94 16.75 15.13 6.42 7.56 7.09 9.77
+D2O 15.53 12.61 17.64 29.0 21.71 8.49 16.56 16.05 16.63 52.5 81.03 20.2 31.88 33.47 26.66
+FastKV 10.55 13.49 13.53 13.27 6.8 4.11 11.94 14.43 11.92 27.0 52.05 17.22 32.1 29.38 18.41
+GemFilter 2.36 3.22 3.99 1.28 2.09 0.78 11.07 13.14 10.9 12.5 36.38 15.54 21.16 26.61 11.50
+SnapKV 11.49 12.47 13.61 12.79 7.05 3.48 11.96 14.61 12.16 28.0 52.11 17.03 32.48 28.39 18.40
+ThinK 8.38 14.5 14.71 20.13 22.21 4.94 11.97 16.44 11.63 18.0 19.59 9.46 21.87 17.09 15.07
+PyramidKV 18.17 23.8 35.97 28.42 31.7 8.76 15.85 16.64 16.2 64.5 83.65 32.41 60.38 54.89 35.10
+FreqKV 19.44 20.2 31.68 35.89 32.0 12.52 15.96 17.39 15.91 60.5 83.98 33.07 62.47 56.57 35.54
+LLaMA-3-instruct-8B
+100% Full Cache 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
+50% LaCache 22.95 28.91 40.99 46.34 36.26 23.02 16.92 16.02 15.59 74.0 91.11 32.97 55.09 49.31 39.25
+D2O 26.61 28.29 38.85 44.95 36.57 22.9 27.66 23.34 25.62 74.0 90.31 41.79 57.57 54.54 42.36
+GemFilter 23.6 29.26 43.71 42.18 36.35 24.48 17.75 16.46 15.7 73.0 90.79 33.3 53.19 38.54 38.45
+FastKV 23.88 29.67 40.75 44.66 34.02 23.07 16.86 16.45 15.61 74.0 90.31 42.02 59.21 53.3 39.63
+SnapKV 22.99 31.89 41.11 43.92 36.42 23.38 17.14 16.45 15.53 74.5 90.23 33.72 59.03 53.97 40.02
+ThinK 21.98 30.97 41.9 45.1 34.18 22.51 27.73 23.38 26.4 74.5 90.16 41.5 62.05 58.37 42.91
+PyramidKV 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
+FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
+1% LaCache 14.27 5.87 20.73 40.86 33.07 18.58 15.01 7.76 10.89 40.0 77.92 16.23 40.7 40.51 27.31
+D2O 24.95 12.77 20.19 36.61 27.2 15.43 22.57 15.18 15.12 56.0 89.34 24.75 32.79 36.18 30.65
+GemFilter 11.8 6.48 19.3 15.23 15.35 7.8 12.28 14.25 9.84 53.42 54.19 20.51 19.37 21.79 20.12
+FastKV 19.57 12.61 31.28 42.59 29.44 20.71 13.96 15.32 12.86 61.0 87.79 29.55 56.74 54.18 34.83
+SnapKV 18.29 14.25 31.87 38.03 26.82 20.03 14.29 14.57 13.08 60.5 88.52 29.95 55.89 53.36 34.25
+ThinK 18.02 10.38 31.46 39.1 26.08 17.46 17.31 21.48 17.97 39.5 88.18 37.27 54.69 55.19 33.86
+PyramidKV 22.79 31.91 41.24 44.2 36.67 22.53 17.29 16.52 15.57 74.5 90.23 33.3 58.95 53.12 39.92
+FreqKV 22.86 39.7 36.44 44.29 39.59 22.57 29.1 22.92 28.35 65.0 89.0 41.9 62.17 55.42 42.81
+4.3 Long Context Understanding
+To further assess the performance on long-prefilling scenarios, we apply FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K.
+Evaluation is conducted on three long context understanding benchmarks: LongBench, RULER, and Needle-in-a-Haystack.
+LongBench.
+Scores on the 5 categories of LongBench are reported in Table 3.
+FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks.
+We compare our method against eviction-based strategies GemFilter (Shi et al., 2024), SnapKV (Li et al., 2024), PyramidKV (Cai.
+et al., 2024b), LaCache (Shi et al., 2025), and FastKV (Jo et al., 2025), merging-based strategies CaM (Zhang et al., 2024), KVMerger (Wang et al., 2024), and D2O (Wan et al., 2025), and a channel compression method ThinK (Xu et al., 2025) that improves SnapKV.
+FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score.
+In particular, it shows significant gains on question answering and summarization tasks such as HotpotQA, 2WikiMQA, and QMSum.
+Moreover, even with an extremely low retaining ratio ($\gamma=1\%$), where other methods degrade sharply, FreqKV remains stable and effective, highlighting its robustness under aggressive compression.
+Table 4: Scores of different KV compression methods on RULER. D2O encounters the OOM issue with the evaluation length of 16K.
+Evaluation Length Method CWE FWE QA VT NIAH Avg.
+8K LLaMA-3 96.40 82.33 62.50 95.00 97.81 86.81
++D2O 93.20 81.00 62.00 88.80 47.53 74.51
++GemFilter 96.40 82.33 62.50 95.00 97.81 86.81
++FastKV 86.10 77.67 63.50 93.00 87.03 81.46
++SnapKV 96.00 80.33 62.50 93.60 92.28 84.94
++ThinK 95.7 78.67 62.00 96.8 94.91 85.62
++PyramidKV 96.40 82.33 62.50 95.00 97.81 86.81
+  +FreqKV 91.80 84.67 64.50 100.00 97.44 87.68
+16K +D2O - - - - - OOM
++GemFilter 0.00 0.00 0.50 0.00 0.00 0.10
++FastKV 0.00 0.00 0.00 0.00 0.00 0.00
++SnapKV 0.00 0.00 0.50 0.00 0.00 0.10
++ThinK 0.00 0.00 0.00 0.00 0.00 0.00
++PyramidKV 0.00 0.00 0.50 0.00 0.00 0.10
+  +FreqKV 45.10 87.33 34.00 44.60 25.47 47.30
+RULER.
+Results on RULER are presented in Table 4.
+FreqKV consistently outperforms baselines across different tasks.
+Notably, baseline methods collapse entirely at 16K evaluation length, which exceeds the original context window of LLaMA-3-instruct (8K).
+This failure arises because they compress the long prompt after the first decoding step.
+Out-of-bound position embeddings are used when calculating attention, resulting in a performance breakdown.
+In contrast, FreqKV remains robust and demonstrates its effectiveness for context extension.
+FreqKV first conducts self-attention within the context window, and then compresses the KV states for out-of-window tokens to get in.
+Compressing key states before applying RoPE, FreqKV extends the context window without requiring position extrapolation or interpolation.
+(a) GemFilter.
+(b) FastKV.
+(c) SnapKV.
+(d) ThinK.
+(e) PyramidKV.
+(f) FreqKV.
+Figure 4: The Needle-in-a-Haystack results of different KV compression methods on LLaMA-3-instruct-8B, with the x-axis representing the document length ranging from 1K to 16K tokens, and the y-axis showing the position of the “needle” within the document.
+Needle-in-a-Haystack.
+Figure 4 illustrates evaluation results on Needle-in-a-Haystack.
+FreqKV performs well in extending the context window of LLaMA-3-instruct.
+In contrast, other KV compression methods fail beyond the original window boundary.
+This breakdown is consistent with our observations in RULER.
+4.4 Long Context Generation
+We use LongGenBench to evaluate the long-context generation performance of FreqKV.
+This benchmark constructs extended generation tasks by concatenating multiple chain-of-thought questions and answers, followed by several real questions in GSM8K and MMLU for long-context responses.
+The model is then required to generate answers with a maximum output length of 4096 tokens.
+Table 5: Performance on LongGenBench.
+Method GSM8K MMLU
+LLaMA-3 42.8 23.8
++D2O 8.0 24.8
++GemFilter 48.5 25.4
++FastKV 48.5 25.4
++SnapKV 49.6 25.8
++ThinK 45.5 25.7
++PyramidKV 49.6 26.2
+  +FreqKV 54.2 31.2
+The accuracy scores (%) are provided in Table 5.
+FreqKV demonstrates substantial improvements in long-generation capability and consistently outperforms existing KV compression methods across both GSM8K and MMLU.
+FreqKV preserves essential contextual information through iterative frequency-domain compression, enabling robust long-context decoding.
 5 Analysis
-Figure 4: Perplexity evaluation on the validation set of PG-19 with different retaining ratios.
-Figure 5: Decoding time with the full cache and FreqKV on the increasing sequence length.
-6.1 Retaining Ratio
-We train LLaMA-2-7b on RedPajama (Computer,, 2023) with different retaining ratios.
-The training length is 8K.
-Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
-The performance of FreqKV with different retaining ratios is presented in Figure 5.
-FreqKV performs better with more frequency components retained.
-Employing an iterative compression manner, less compressions are performed on the more recent tokens by FreqKV, which benefits long context language modeling.
-It consistently achieves low perplexity as the evaluation context length grows from 2K to 64K, far exceeding the training length.
-6.2 Decoding Latency
-Table 3: The decoding latency (seconds) of FreqKV and Dropping on LLaMA-2-7b with the increasing sequence length.
-The difference between FreqKV and Dropping shows the compression latency introduced by FreqKV.
-Models 4K 8K 16K 24K 32K
-Compression Times 0 3 7 11 15
-Dropping 123.78$\pm$1.46 253.32$\pm$5.76 506.60$\pm$5.11 760.46$\pm$7.15 1022.50$\pm$18.83
-FreqKV 124.34$\pm$1.80 252.75$\pm$3.55 507.61$\pm$5.44 764.82$\pm$9.72 1029.00$\pm$19.97
-FreqKV performs compressions on-the-fly during decoding.
-We compare the decoding time required by LLaMA-2-7b with the full cache and our FreqKV when the sequence length increases.
-As shown in Figure 5, the decoding time starts to diverge at the length of 4K.
-While the full cache utilization leads to a quadratic growth in decoding time, the decoding time of FreqKV increases approximately linearly with a negligible time spent on compression, showcasing its efficiency.
-Furthermore, we measure the decoding latency of FreqKV and Dropping with the increasing length.
-Dropping removes the earliest tokens instead of compression in the frequency domain.
-The difference in the decoding time between the two methods shows the latency of compression.
-The sink size is set to 4 and the retaining ratio is 0.5.
-Table 3 reports the average decoding time (seconds) over five runs, along with standard deviations.
-It can be observed that the additional latency (6.5s) introduced by FreqKV accounts for only 0.64% of the total decoding time at 32K tokens.
-Further analysis for the compression overhead is provided in Appendix D.
+5.1 Efficient Extrapolation
+Table 6: Extrapolation performance of FreqKV on PG-19 up to 256K context length.
+We measure latencies (seconds) and PPL scores at different lengths.
+“Compression Count” denotes the number of compressions during decoding.
+“Compression Latency” stands for the latency introduced by all the frequency-domain compressions during decoding.
+Evaluation Length 8K 16K 32K 64K 128K 256K
+Compression Count 3 7 15 31 63 127
+Decoding Latency 214.96 445.69 907.49 1832.77 3692.03 7422.64
+Compression Latency 0.62 (0.29%) 1.09 (0.24%) 2.18 (0.24%) 4.00 (0.22%) 8.31 (0.23%) 16.39 (0.22%)
+PPL 7.76 7.74 7.73 7.73 7.73 7.73
+Furthermore, we evaluate the performance of FreqKV at longer lengths on the validation set of PG-19 as shown in Table 6.
+With the training length limited to 8K tokens, it effectively extrapolates the context window of LLaMA-2-7B to 256K, which is 64 times larger than the original size.
+This is enabled by its iterative compression strategy, which applies fewer compressions to recent tokens while progressively compressing earlier ones.
+Consequently, FreqKV maintains low perplexity as the evaluation context length scales from 8K to 256K, far exceeding the training length.
+In addition, latencies (seconds) of decoding and frequency-domain compression at different lengths are also reported in Table 6.
+It shows that the decoding time of FreqKV increases approximately linearly, with a negligible time spent on compression.
+The overhead of the compression accounts for less than 0.3% and does not grow with the context length since the number of compressions increases linearly with the context length.
+Further analysis for the compression overhead is provided in Appendix G.
+In summary, FreqKV demonstrates strong extrapolation ability to ultra-long contexts while preserving both efficiency and stability.
+5.2 Training Cost
+The training time and memory requirements of FreqKV and Full FT on LLaMA-2-7B are reported in Table 7.
+8 NVIDIA RTX6000 Ada-48GB GPUs are used.
+For comparison, we set the batch size per device to 1 and gradient accumulation steps to 8.
+Statistics show that FreqKV significantly reduces both training time (by 40.89%) and memory requirements (by 44.19%), demonstrating its computational efficiency.
+Table 7: The training time and memory requirements of FreqKV and Full FT.
+Method Training Length Time (hours) Memory (GB)
+Full FT 8K 28.88 44.04
+FreqKV 8K 17.07 24.58
+16K 38.57 34.00
+32K 89.35 45.37
 6 Conclusion
-In this paper, we introduce FreqKV to compress KV states iteratively in the frequency domain for LLMs.
-We exploit the energy concentration of KV states in the frequency domain within the decoder layers.
-Specifically, we filter out the high-frequency components that are of low magnitude and retain the low-frequency components for compression.
-The KV cache is compressed in the frequency domain without introducing additional compression modules.
-Iteratively compressing the KV cache, FreqKV could extend the context window efficiently for LLMs.
-With minimal training of low-rank adaption, LLMs learn to leverage the compressed KV cache.
-Through extensive experiments and analysis on long context modeling and understanding, FreqKV demonstrates its efficiency and effectiveness in context extension.
-Appendix A Head-wise Analysis of Power Spectrums
-We explore the power spectrum distribution of key states and value states in different attention heads of LLaMA-2-7b as shown in Figure 6.
-Although values of power spectrums vary in different heads, their distribution exhibits similar patterns.
-They have a consistent tendency to aggregate energy in the low-frequency components along the decoding procedure.
-It could be promising to study specific differences and associations in different heads or other modules.
+In this paper, we introduce FreqKV, a frequency-domain KV cache compression framework that exploits the energy concentration of KV states to retain low-frequency components while discarding redundant high-frequency ones for LLMs.
+By iteratively compressing the KV cache, FreqKV achieves efficient long-context extension while preserving decoding quality.
+Without introducing additional parameters or architectural modifications, FreqKV exhibits remarkable efficiency and effectiveness on both prefilling and decoding stages.
+Extensive experiments across language modeling, generation, and understanding benchmarks demonstrate that FreqKV not only consistently surpasses existing KV compression methods but also offers a viable and efficient approach for context window extension.
+Ethics Statements
+Our work pertains to key-value compression and context extension of large language models.
+In this work, we use only publicly available data and artifacts.
+There are no ethical issues in our paper, including its motivation and experiments.
+Reproducibility Statements
+We have provided detailed implementations of our method throughout the paper.
+Our method is comprehensively elaborated in Section 3.
+Detailed settings of our experiments and analyses are given in Section 4,  5, and Appendix F.
+Our code is included in the Supplementary Material for reproducibility.
+The Use of Large Language Models
+We used ChatGPT solely as a language editing tool to polish grammar, improve clarity, and refine the academic style of the manuscript.
+All research ideas, methods, experiments, and analyses were independently developed and conducted by the authors.
+Appendix A Discrete Cosine Transform
+The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain.
+Several variants of DCT exist, with DCT-II being the most common.
+For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
+[…]
+where $t=0,1,\cdots,N-1$.
+$\alpha_{t}$ is the normalization factor.
+The time-domain signal ${\bm{X}}_{0:N-1}$ can be recovered by applying the inverse DCT (IDCT) on the frequency components ${\bm{Z}}_{0:N-1}$:
+$x_{n}=\sum_{t=0}^{N-1}\alpha_{t}\cdot z_{t}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right].$ (7)
+The frequency components are expressed as a combination of the original signals.
+The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$.
+The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components.
+The components of higher energy in the frequency domain indicate they are more informative (He et al., 2023).
+For time signals (e.g., texts) and space signals (e.g., images), their energy is most concentrated in low-frequency DCT components.
+This phenomenon occurs due to the mathematical properties of DCT and the inherent smoothness of real-world signals.
+The DCT expresses a signal as a sum of cosine functions oscillating at different frequencies.
+In the case of text embeddings, neighboring tokens exhibit strong temporal correlations, meaning adjacent values in the semantic space tend to be similar.
+Low-frequency cosine components are slowly varying and capture these smooth variations well, while high-frequency components mainly represent abrupt changes or noise.
 (a) The average power spectrums of key states.
@@
 Figure 5: The average power spectrums of key states and value states in different heads of Llama-2-7b.
-Appendix B Compressing Key States with RoPE
+Appendix B Head-wise Analysis of Power Spectrums
+We explore the power spectrum distribution of key states and value states in different attention heads of LLaMA-2-7B, as shown in Figure 5.
+Although the values of power spectrums vary in different heads, their distribution exhibits similar patterns.
+They have a consistent tendency to aggregate energy in the low-frequency components along the decoding procedure.
+It could be promising to study specific differences and associations in different heads or other modules.
+Appendix C Compressing Key States with RoPE
 Figure 6: The average power spectrums of key states with RoPE.
@@
 In contrast, FreqKV compresses and caches key states before applying RoPE.
-Therefore, they are allocated with position indices within the original window rather than the lengthy sequence, which results in the context window extension with no need of using out-of-bound position embeddings.
+Therefore, they are allocated with position indices within the original window rather than the lengthy sequence, which enables context window extension with no need of using out-of-bound position embeddings.
 Table 8: Perplexity evaluation on the test sets of PG-19 and Proof-pile. “Variant” stands for FreqKV with key states compressed after applying RoPE.
@@
 Variant 3.16 2.88 2.80
-Appendix C Training settings
-For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer,, 2023) pre-training dataset for 1000 steps.
-For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al.,, 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
+Appendix D Training settings
+For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama (Computer, 2023) pre-training dataset for 1000 steps.
+For long context understanding, we SFT LLaMA-2-chat and LLaMA-3-instruct on the instruction following dataset LongAlpaca (Chen et al., 2024) for 5 epochs, which consists of 6.28K long-context QA samples.
 The total batch size ($GPU\_number\times Batch\_size\_per\_device\times Gradient\_accumulation\_steps$) is 64.
 The learning rate increases linearly from 1e-6 to 2e-5 with 20 warm-up steps and remains constant in the following steps.
-The rank used in the LoRA (Hu et al.,, 2021) fine-tuning is set to 8.
-Following LongLoRA (Chen et al.,, 2024), the embedding and normalization layers are learnable during training.
-All the experiments are conducted on ADA6000 and H100 GPUs.
-Moreover, we equip our method with FlashAttention-2 (Dao,, 2023) for further acceleration and memory saving.
-The training time and memory requirements of FreqKV and Full FT on the RedPajama are reported in Table 5.
-For comparison, we set $Batch\_size\_per\_device$ to 1 and $Gradient\_accumulation\_steps$ to 8.
-8$\times$×ADA6000-48GB GPUs are used.
-Statistics show that FreqKV significantly reduces both training time and memory requirements, demonstrating its efficiency.
-Table 5: The training time and memory requirements of FreqKV and Full FT.
-Methods Training Length Time (hours) Memory (GB)
-Full FT 8K 28.88 44.04
-FreqKV 8K 17.07 24.58
-16K 38.57 34.00
-32K 89.35 45.37
-Appendix D Compression Overhead
+The rank used in the LoRA (Hu et al., 2021) fine-tuning is set to 8.
+Following LongLoRA (Chen et al., 2024), the embedding and normalization layers are learnable during training.
+Moreover, we equip our method with FlashAttention-2 (Dao, 2023) for further acceleration and memory saving.
+Table 9: Perplexity evaluation on the test set of Proof-pile.
+The superscript “*” means that we reproduce LoCoCo following their official code for evaluation.
+The results of full fine-tuning and LongLoRA are reported from LongLoRA (Chen et al., 2024).
+Full FT and LoCoCo encounter the OOM (Out-of-Memory) issue with the training length of 16K.
+Size Training Length Method Inference Cache Evaluation Context Length
+2048 4096 8192 16384 32768
+7B 8192 Full FT Full 3.14 2.85 2.66 - -
+LoCoCo Compressed 3.40 3.20 2.88 - -
+LongLoRA Full 3.20 2.91 2.72 - -
+FreqKV Compressed 3.15 2.88 2.79 2.76 2.75
+16384 Full FT Full OOM during Training
+LoCoCo Compressed OOM during Training
+LongLoRA Full 3.17 2.87 2.66 2.51 -
+FreqKV Compressed 3.15 2.88 2.78 2.74 2.73
+32768 LongLoRA Full 3.35 3.01 2.78 2.61 2.50
+FreqKV Compressed 3.16 2.89 2.78 2.74 2.72
+13B 8192 Full FT Full 2.96 2.69 2.53 - -
+LongLoRA Full 3.04 2.77 2.60 - -
+FreqKV Compressed 2.99 2.74 2.65 2.63 2.63
+Appendix E Evaluation on Proof-pile
+In Table 9, we present the evaluation results on the Arxiv math dataset Proof-pile.
+On Proof-pile, FreqKV trails LongLoRA at longer context lengths.
+This gap arises because Proof-pile, unlike natural language corpora, contains dense symbolic and mathematical expressions where minor variations (e.g., a single symbol or index) drastically alter semantics.
+Such localized variations are more difficult to preserve under compression, making it particularly challenging for models using compressed cache to predict the next token.
+LongLoRA retains the full KV cache during inference, avoiding any information loss and thus performing better on formula-heavy tasks.
+Nonetheless, FreqKV remains competitive and notably outperforms another compression-based method, LoCoCo.
+Additionally, the performance of FreqKV does not degrade further as the context length increases, highlighting its robustness.
+Appendix F Ablation Study
+F.1 Retaining High/Low-Frequency Components
+We further compare variants of FreqKV that retain only high-frequency vs. only low-frequency components using LLaMA-2-base-7B on the validation set of PG-19. Results (PPL) are provided in Table 10.
+It shows that low-frequency components adapt better than the variant that retains high-frequency parts, which is consistent with our analysis in summarization tasks (Table 1).
+While the variant deteriorates when exceeding the original context window, low-frequency components, which carry more information, benefit the model, and the perplexity remains stable.
+This result aligns well with Frequency Principle (Xu et al., 2019; Luo et al., 2019) that low frequencies converge faster than high frequencies for Deep Neural Networks.
+Table 10: Perplexity evaluation on PG-19 with high/low-frequency components retained.
+Components Retained Evaluation Context Length
+2048 4096 8192 16384
+high-frequency 8.29 7.92 8.42 8.70
+low-frequency 8.20 7.84 7.74 7.73
+Figure 7: Perplexity (Line) and decoding latency (Bar) with different retaining ratios.
+Figure 8: Decoding time with the full cache and FreqKV on the increasing sequence length.
+F.2 Retaining Ratio
+We train LLaMA-2-7B on RedPajama (Computer, 2023) with different retaining ratios.
+The training length is 8K.
+Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
+The performance of FreqKV with different retaining ratios is presented in Figure 8.
+The decoding time (seconds) is averaged over five runs.
+FreqKV performs better with more frequency components retained.
+However, larger retaining ratios lead to more frequent compression steps and denser attention computation, hence higher latency.
+We also evaluate the performance of FreqKV when modifying its retaining ratio during evaluation on the validation set of PG-19.
+The PPL scores at a length of 8K are presented in Table 11.
+These results suggest that FreqKV is robust to mismatched training/evaluation retaining ratios, especially when using larger inference ratios than training (which retains more information).
+Table 11: Perplexity evaluation on PG-19 with mismatched training/evaluation retaining ratios.
+Retaining Ratio during Training Retaining Ratio during Evaluation
+0.25 0.5 0.75
+0.25 7.84 7.81 7.82
+0.5 7.89 7.76 7.77
+0.75 7.97 7.79 7.75
+F.3 Sink Size
+The effect of sink token count has been explored in StreamingLLM (Xiao et al., 2024b).
+We also ablate the effect of the sink token number $S$ for FreqKV on LLaMA-2-7B.
+Results are given in Table 12.
+It shows that a threshold of four sink tokens appears to be enough, with subsequent additions contributing marginal effects.
+Table 12: Perplexity evaluation on PG-19 with different numbers of sink tokens.
+Sink Size Evaluation Context Length
+2048 4096 8192
+0 8.23 7.87 7.84
+2 8.21 7.85 7.78
+4 8.20 7.84 7.76
+8 8.20 7.83 7.75
+F.4 Effect of Frequency-Domain Compression
+(a) Dropping.
+(b) FreqKV.
+Figure 9: Comparison of FreqKV and its variant Dropping on LLaMA-2-chat-7B.
+To demonstrate the effectiveness of our compression operation, a variant method, Dropping, is also implemented and evaluated on Needle-in-a-Haystack for comparison, as in Figure 9.
+Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al., 2024b).
+Although trained under identical conditions, Dropping suffers from an irreversible loss of early-context information.
+FreqKV, by contrast, preserves essential signals and achieves superior performance.
+F.5 Effect of rescaling
+Figure 10: Curves of training loss for FreqKV with and without rescaling.
+In Equation 3, $\sqrt{\frac{L}{N}}=\sqrt{\gamma}$ works as a rescaling factor for the compressed signals to restore the original amplitude when conducting IDCT.
+In DCT, an $N$-length time-domain signal is normalized by $\sqrt{N}$; during IDCT, the retained $L$ frequency components are correspondingly normalized by $\sqrt{L}$.
+When iterative compression is applied across multiple steps, this mismatch in normalization scales can cause the reconstructed signals to be progressively amplified.
+In our pilot study, the training loss is significantly higher at the early stages when the compressed signals are not rescaled as presented in Figure 10.
+Therefore, we introduce the rescaling factor $\sqrt{\frac{L}{N}}$ to restore the original amplitude in the time domain.
+Appendix G Compression Overhead
 We conduct further studies on the compression overhead of FreqKV.
+FreqKV performs compressions on-the-fly during decoding.
+We compare the decoding time required by LLaMA-2-7B with the full cache and our FreqKV when the sequence length increases.
+As shown in Figure 8, the decoding time starts to diverge at the length of 4K.
+While the full cache utilization leads to a quadratic growth in decoding time, the decoding time of FreqKV increases approximately linearly, with a negligible time spent on compression, showcasing its efficiency.
 As introduced in Section 3.4, chunk-wise attention is performed for the prefilling tokens.
 The size of the attention matrix in each chunk is $(N-L-S)\cdot N$ except for the last few tokens.
-Therefore, the computational cost of self-attention grows approximately linearly with the input length like sliding window attention (Beltagy et al.,, 2020).
+Therefore, the computational cost of self-attention grows approximately linearly with the input length, like sliding window attention (Beltagy et al., 2020).
 We use torchprofile [^1] https://github.com/zhijian-liu/torchprofile to count the number of Floating Point Operations (FLOPs) with input sequences of different lengths for LLaMA-2-7B.
-We calculate FLOPs for the model with the original attention which leverages full KV states, and with FreqKV with the retaining ratio of 0.5.
-To quantify the compression overhead, we have measured FLOPs of the Dropping, which retains the most recent tokens instead of compressing in the frequency domain.
+We calculate FLOPs for the model with the original attention, which leverages full KV states, and with FreqKV with the retaining ratio of 0.5.
+To quantify the compression overhead, we have measured FLOPs of Dropping, which retains the most recent tokens instead of compressing in the frequency domain.
 It shares the same sink size and retaining size as FreqKV.
@@
 FreqKV reduces more FLOPs as the input length grows from 4K to 32K compared to Full KV.
-This is because the compression is performed every $N-L-S$ tokens with the complexity of $O(N\text{log}N)$, which is negligible compared to the quadratic self-attention.
+This is because the compression is performed every NLSN-L-S tokens with the complexity of $O(N\text{log}N)$, which is negligible compared to the quadratic self-attention.
 Table 13: FLOPs (TFLOPs) with input sequences of different lengths.
-Experiments are conducted on the ADA6000 GPU of 48GB, where full KV of 12K tokens with float16 will cause an OOM (Out-of-Memory) issue.
+Experiments are conducted on a single NVIDIA RTX6000 Ada-48GB GPU, where full KV cache of 12K tokens with float16 will cause an OOM (Out-of-Memory) issue.
 The difference between FreqKV and Dropping shows the computation overhead of compression.
 Models 4K 8K 12K 16K 32K
-Full KV 62.93 143.46 OOM OOM OOM
+Full Cache 62.93 143.46 OOM OOM OOM
 Dropping 62.93 125.86 188.79 251.72 503.43
@@
 Compression Overhead 0 (0%) 0.039 (0.031%) 0.064 (0.034%) 0.090 (0.036%) 0.193 (0.038%)
-Compression Times 0 3 5 7 15
+Compression Count 0 3 5 7 15